OpenStudy (anonymous):

Algebraic Fractions. Oh no! [(2a^2-9a-35)/(12x^3)]x[(9x^2)/(2a^2-3a-20)] I got the answer of... [3(a-7)]/[4(a-4)x] But it says that it is wrong.. Helpp! Thank you!

OpenStudy (anonymous):

I don't know how to use the equation thing to make it look normal, so I hope you can understand! Sorry!

mathslover (mathslover):

\[\large{\frac{2a^2-9a-35}{12x^3}\times x \times \frac{9x^2}{2a^2-3a-20}}\] is this the question?

OpenStudy (anonymous):

There is no x in the middle, it is just multiplying the two fractions. Sorry if that was confusing..

mathslover (mathslover):

ok no problem : \[\large{\frac{2a^2-9a-35}{12x^3} \times \frac{9x^2}{2a^2-3a-20}}\] \[\large{\frac{2a^2+14a-5a-35}{12\cancel{x^3}^x}\times \frac{9\cancel{x^2}^1}{2a^2-3a-20}}\]

OpenStudy (anonymous):

Wouldn't +14a-5a give you +9 ?a

OpenStudy (anonymous):

**+9a

mathslover (mathslover):

very right my mistake sorry : \[\large{\frac{2a^2-14a+5a-35}{12\cancel{x^3}^x}\times \frac{9\cancel{x^2}^1}{2a^2-3a-20}}\]

OpenStudy (anonymous):

Okay, just checking !

mathslover (mathslover):

\[\large{\frac{2a^2-14a+5a-35}{12\cancel{x^3}^x}\times \frac{9\cancel{x^2}^1}{2a^2-3a-20}}\] \[\large{\frac{2(a-7)+5(a-7)}{12x}\times \frac{9}{2a^2-8a+5a-20}}\] \[\large{\frac{(2a+5)(a-7)}{12x}\times \frac{9}{(2a+5)(a-4)}}\]

mathslover (mathslover):

\[\large{\frac{\cancel{2a+5}^1(a-7)}{12x}\times \frac{9}{\cancel{2a+5}^1(a-4)}}\] \[\large{\frac{(a-7)}{12x}\times \frac{9}{a-4}}\]

OpenStudy (anonymous):

Ohh. Okay. & One question.. Wouldn't it be 2(a^2-7) ?

OpenStudy (anonymous):

In your second step on the left numerator

OpenStudy (anonymous):

Well I guess technically, your third.

mathslover (mathslover):

no i meant : it was : 2a(a-7)

OpenStudy (anonymous):

Oh, right ! Oh, well thank you. You made it super easy to see when I originally made my mistake :) Thank you soo much ! (:

mathslover (mathslover):

welcome @theequestrian and sorry for the mistakes made byy me

OpenStudy (anonymous):

Oh no problem!!